[Raymond, Aidan, Chris, Koji]
P6 element (500um)^2
- We looked at the current amp (FEMTO) output. The amplifier saturated at the gain of 10^3 V/A. Looking at the output with a scope, we found that there is a huge 1.2MHz oscillation. Initially, we thought it is the amplifier oscillation. However, this oscillation is independent of the amplifier bandwidth when we tried the our-own made transimpedance amp.
- Shorting the cryostat chamber to the optical table made the 1.2MHz significantly reduced. Also, connecting the shield of the TEC/Laser controller made the oscillation almost invisible. This improvement allowed us to increase the amp-gain up to 10^7.
- Then the dominant RMS was 60Hz line. This was reduced by more grounding of the cable shields. The output was still dominated by the 60Hz line, but the gain could be increased to 10^8. This was sufficient for us to proceed to the careful measurements.
- The dark current was measured by the source meter, while the photocurrent (together with the dark current) was measured under the illumination of the ~1mW light on the PD.
- Attachment 1 shows the dependence of the dark current against the swept bias voltage. We had ~mA dark current at the room temp. So, this is ~10^5 improvement.
- Attachment 2 shows the dependence of the apparent QE against the swept bias voltage. The dark current was subtracted from the total current, to estimate the contribution of the photocurrent in the measurement.
- Attachment 3 shows the dark noise measurement at the reverse bias of ~0.6V. Up to 1kHz, the noise level was below the equivalent shotnoise level of 1mA photocurrent.
All the data and python notebook in the attached zip file.
We placed a power meter after the fiber collimator, 75mm focal length lens and HR mirror at 45 degrees - basically, we placed the power meter immediately before the input window to the cryo chamber after all the intervening optics from the fiber output.
For a series of laser diode current levels, we measured the power on the power meter and the corresponding voltage on the reference photodetector that is monitoring a 10% pick-off from the laser. The calibration is as follows:
Wow. This is great, thanks Chris.
Note: in the preceding table, channel numbers use the digital convention (numbered starting from zero), which is not the convention used by the AA/AI chassis front panel (numbered starting from one).
I can see some screws are not vented. You also need to use a vented screw for the additional temp sensor if the face screws of the PD mounts are not vented.
You can use a bunch of clean clamps and screws I brought. They are in a mylar bag.
If you need more vented screws, please specify the size and length. I can grab some from the 40m cleanroom.
The IR Labs cryostat has its internals wired and attached to the baseplate. PD A2 was clamped and the vacuum pumps turned on for the first cooling test.
[in the morning I will update with a detailed pin-out and label the attached photo (labeled 12/13, pin out in separate post)]
Recently Duo wanted to make an arbitrary waveform excitation using the QIL cymac, but it wasn't working. An excitation would die after 10 seconds or so, with awgtpman reporting that the data was too far in the future.
It turns out this was caused by a missing leap second in the RTS software. It is now fixed upstream, and we're running a patched version of awgtpman on fb4, until the change propagates to the packaged version.
I borrowed KEITHLEY 2450 source meter from Rich. The unit comes with special coaxial cables and banana clips. Most of the peripherals are evacuated in the OMC lab.
The dark current of A2P2, A2P3, A2P6 were measure with different temperatures (300K, 270K, 254K). The plot combined with the previous measurement ELOG QIL 2425.
== How to use the source meter ==
- Two-wire mode: Connect the wires to the diode
- Over voltage protection: [MENU] button -> SOURCE / SETTINGS->Over Voltage Protectiuon 2V
- Sweep setting: [MENU] button -> SOURCE / SWEEP -> e.g. Start -750mV, Stop +500mV, Step 10mV, Source Limit 1mA -> Select Generate
- Graph View: [MENU] button -> VIEWS / GRAPH
- Start measurement: Note: The response of [TRIGGER] button is not good. You need to push hard
This starts the sweep, or a menu shows up if your push is too long -> Select "Initiate ..."
- Data Saving: [MENU] button -> MEASURE / READING BUFFERS -> Save to a USB stick
Update on the 32 pin female connector:
Mouser's overnight delivery was rejected on 12/3 by mail services for being soaked in an "unkown liquid" and was therefore taken away by the courier for return to Mouser. This was the last one in stock, so I ordered a replacement through Digikey for express delivery this morning 12/4, but it has not yet arrived. I've called Fedex and the package was sorted at the LA facility but not given to the courier for morning delivery. It is now estimated for delivery this afternoon.
The quantities we want to measure as a function of the temperature:
- Temperature: 2.2k thermister resistance / 100ohm platinum RTD
- QE (Illuminating output / Dark output / Reference voltage / Reference dark output)
- Dark current (vs V_bias) -> Manual measurement or use a source meter
- Dark noise (PSD) 100kHz, 12.8k, 1.6kHz, 100Hz
[Raymond and Koji]
We dunked the PD socket test piece into LN2 and repeated heat cycle 8 times. No obvious change was observed. Then the wires were pulled to find any broken joint or etc.
None of the solder joints showed the sign of failure.
For cleanliness, we are going to use In-Ag solder (no flux) for the actual wiring.
Attachment 1: Frozen connector
Attachment 2-4: Inspection after thawing.
While I'm still waiting for the proper connector for the vacuum feedthru of the IRLabs cryostat, I have connected to the Dsub9/15 split cable to another Dsub9 connector so that I can test the cooling of the InAsSb sensor in air. Also, the 2004nm laser, a fiber-coupled faraday isolator, and 90:10 beam splitter was moved to the cryostat table and fixed on a black al breadboard. [Attachment 1]
The InAsSb TEC was controlled by the TEC controller of ITC-50. I didn't change the PID parameters of the controller but the temperature nicely setteled to the setpoint. The sensor has a 2.2kOhm thermister. And the max current for the TEC was unknown. The TEC driver had the current limiter of 0.3A and it was not changed for now. With this current limit, the thermistor resistance of 10Kohm was realized. This corresponds to the temperature of about -20degC. According to the data sheet given by Alex, the resistance/temperature conversion is given by the formula
1/T = 7.755e-4 + 3.425e-4*log(R)+1.611e-13*log(R)^3
To satisfy the curiosity, the dark current of a (500um)^2 element was measured between -250K and -300K. At -254K, the dark current went down to the level of 40uA (1/15 of the one at the room temp). For the measurement, the bias voltage was set to be 0.5 and 0.6V. However, it was dependent on the diode current. (Probably the bias circuit has the output impedance). This should be replaced by something else.
The external Dsub cable is ready except for the 32pin connector to be plugged-in to the chamber. See QIL ELOG 2460 for the pin assignment.
Normal solder (Sn63 Pb37): with flux, wetting o
Pure Indium - In 99.995: no flux, wetting x, low melting temp, like paste
Pb93.5 Sn5 Ag1.5: with flux, wetting o, high melting temp (soldering iron setting 380~430F)
Cryo solder In97 Ag3: no flux, wetting x, low melting temp, like paste
The IR Labs cryostat is now pumping down. The thor labs posts are necessary for centering the optical port at a height of 5.5". The orange cabinet used to elevate the pumping station was relocated from next to the computer (was, and still is, empty). The power supply is for the wide range vacuum gauge attached to the pumping line.
I've ordered a liquid nitrogen dewar for arrival Monday morning.
For future reference, the gauge communicates via a VGA port, for which the pin-to-wire association is as such: 1-Black; 2-Brown; 3-Red; 4-Orange; 5-Yellow; 6-Green; 7-Blue; 8-Purple; 9-Grey; 10-White; 11-Pink; 12-Cyan; 13-Black/White; 14-Brown/White; 15-Red/White (MKS 901P non-ethercat wiring diagram)
edit by RXA: replaced multi GB PDF w a reasonable JPG.
System diagram of the PD QE test with the IRLabs cryostat.
PT-SE (MS/PT-SE) connector data sheets
Amphenol catalog http://www.amphenol-industrial.com/images/catalogs/PT.pdf
Detoronics Hermeic Sealed Connectors (DT02H-18-*PN) http://www.hselectronics.com/pdf/Detoronics-Hermetic-Connectors.pdf
AF8 crimping tool (expensive!) https://www.mouser.com/ProductDetail/DMC-Tools/AF8?qs=gvhpkjpQEVSjrLbsepewjg%3D%3D
AF8 alternative https://www.jrdtools.com/?gclid=Cj0KCQiA2vjuBRCqARIsAJL5a-IQ9ztCEYKdo645v_RhUBJS3eMIars1LubjlKZoorS-lnx6ClDDiMUaAlZiEALw_wcB
Thermistor link: https://www.tec-microsystems.com//Download/Docs/Thermistors/TB04-222%205%25%20Thermistor_Specification_upd2018.pdf
TEC spec: Mounted TEC type: 2MD04-022-08/1 https://www.tec-microsystems.com/products/thermoelectric-coolers/2md04-series-thermoelectric-coolers.html
2MD04-022-08/1 dTmax = 96, Qmax = 0.4W, Imax = 0.7A, Umax = 2.0, ACR = 2.29 Ohm
The PD mounts were delivered from ProtoLabs. The order was sent on Tue last week and it's here on Monday. Excellent!
And the quality looks pretty good.
The surfaces are sandblasted. Do we want to do any process on the bottom surface to reduce the thermal resistance?
An indium solder string also came in.
The power transmission of the optical window for the IRLab cryostat was measured to be 0.966+/-0.002 at 2004nm. (Attachment 1)
A chopper powermeter was set to the QE measurement setup (Attachment 2). The window was held with a mount as shown in Attachmnent 3. The laser source was excited with the pumping current of 101.04mA. The output power was monitored with a Thorlabs DET10D (PD#2 with Amp#2) attached at the 10% side of the 90:10 beamsplitter. The detected photocurrent after subtracting the dark current of 15.7uA was 152uA. The power meter detected the power around 0.95mW, while the power with the window inserted was around 0.91~0.92.
PD1 Window No Window
[V] [mW] [mW]
-0.855 0.913 0.944
-0.855 0.906 0.951
-0.855 0.914 0.947
-0.855 0.922 0.950
-0.855 0.913 0.949
-0.855 0.912 0.948
-0.855 0.920 0.946
-0.855 0.915 0.946
-0.855 0.916 0.951
-0.855 0.915 0.952
-0.855 0.919 0.947
-0.855 0.921 0.944
-0.855 0.916 0.948
Note: PD1 had the dark output of -0.0809V.
Note2: The power meter readings had the fluctuation of +/-0.005 mW
PD1 Window No Window
[V] [mW] [mW]
-0.855 0.913 0.944
-0.855 0.906 0.951
-0.855 0.914 0.947
-0.855 0.922 0.950
-0.855 0.913 0.949
-0.855 0.912 0.948
-0.855 0.920 0.946
-0.855 0.915 0.946
-0.855 0.916 0.951
-0.855 0.915 0.952
-0.855 0.919 0.947
-0.855 0.921 0.944
-0.855 0.916 0.948
Salvaged an assort of vented/non-vented screws, washers, spring washers, and clamps for #4-40 & 1/4-20 from the 40m cleanroom stock. They are clean enough for the cryostat use.
Removed the IR Labs cryostat window for testing. It's on the cryocooler bench, see attached photos
Update of our available electrical feedthroughs:
2 x 19 pin round with corresponding internal connectors and external connectors
2 x 15 D sub
- re-tapped the PEEK support pieces to 1/4-20 to match the bottom of the tank, added them with vented socket set screws. These will need to be replaced with brass or aluminum to better match thermal contraction (steel screws likely to crack the PEEK upon cooling)
- Drilled holes (attached, highlighted in orange) in the pump station base in order to utilize the more robust shock absorbers that came with the oil pump.
- Noticed significant flaking of the nuts/bolts when removing the tank lid. Suggested using anit-seize compound on the tank lid bolts but Chris advised against anything lube-like on the system. Just a note to remember to check the integrity of the bolts going forward before tightening the lid to avoid bolt seizing.
Ordered the electrical feedthrough from IR Labs on overnight shipping. I've attached the diagram of the feedthrough, 32 pin is the only available configuration for this size feedthrough.
Will attempt to use my makeshift 12 pin feedthrough if the proper one does not arrive tomorrow (thanks Koji for the 40m drill bit tip). I'll clear the bench of the IR Labs carcasses tomorrow.
UPDATE: feedthrough will arrive by 10:30 on Monday the 25th
Here you are.
The large cryostat has 1/4-20 holes on a 2 inch grid
The IR labs cryostat has 4-40 holes on a 1 cm grid
I'll check the 40m for a bit (5/8-3/4")
Ah, I have designed the PD holder with the venting targeted for 1/4-20 holes with 1" grid...
The dog clamps to hold the PD units also need to be compatible with 4-40 screws.
How big the hole diameter should be? Can you find a suitable drill at the 40m?
In the QIL and noticed that the vent above the cryo chamber is heating the lab while the vent above the 2um laser bench is blasting AC. Both have been running continuously since I first entered the lab (~10:30)
The cold plate breadboard holes are indeed 4-40.
I've installed a new o-ring and vacuum valve, but I've scoured the labs for any piece that fits the open port on the side of the cryostat and have come up empty-handed. I've attached a photo with all of the IR labs input pieces found thus far (except duplicates or ones that match these same dimensions), none fit the drill pattern or counter sunk hole of the cryostat port. I tried but could not find mention of this specific port's removal in the elog history.
Spent a while trying to make a Franken-stat from the innards and limbs of the other decommisioned IR lab cryostats but nothing allowed for a correct optical path. So I then started to make a custom port with a 122 o-ring and an old ISO blank flange, but I need a drill bit larger than 0.5" for the center electrical feedthrough and am having ID card issues accessing the student shop over in CES.
I will call IR labs tomorrow and see if they have an a standard piece for this feedthrough.
Here are the commands.
sudo /sbin/rmmod c4tst c4iop && cd /opt/rtcds/caltech/c4/target/c4iop/scripts/
./startupC4rt && cd ../../c4tst/scripts/ && ./startupC4rt && systemctl start firstname.lastname@example.org && systemctl start email@example.com && systemctl restart firstname.lastname@example.org
See attached photos for the internal layout of Zach's cantilever cryostat (all internal components left in cryo lab (except the heater)) The holes are on a 1cm grid and, I believe, are threaded for 4-40 (will check Monday). The internal 'window' is about an inch off the colplate. There is ~2 inches of space from the colplate to the first radiation shield lid.
The external optical port on the vacuum tank is 1.75" from the bench and is 0.75" in diameter. The cryostat is missing the lid o-ring, electrical feedthrough vacuum port, and intermediary valve between the vacuum space and the vacuum pumps. I will search for these on Monday.
Last week I moved the upper portion of the crane to the new, bolted crane support stand. Chub removed the wheeled lower section from the lab shortly thereafter. I also re-threaded the nylon lifting strap to remove slack and level the lid a bit better during lifting and moved one of the side tables next to the crane so the lid can be safely lowered after being lifted off the tank (see first photo).
Opened the tank today to check internal dimensions. It is now closed (top bolts finger tight) but not under vacuum. The diaphragm pump was dispatched today, so will replace the dirty pump and pull vacuum again upon arrival.
Attached a photo of the baseplate for future drill pattern reference. Note there are three anomalous holes, this is where the PEEK support poles should go. It was discovered today that these holes are 1/4-20 tapped but the PEEK pillars are dirlled/threaded for a smaller bolt.
I've attached a rough cartoon of the cold plate height relative to the optical ports and the tank wall. The outer rad shield is not shown and is slightly misaligned, but it can be easily aligned with a ~1.5 mm shim (better for thermal isolation anyways).
The QE of the (500um)^2 element has been tested with a half-power (0.51mW) instead of 0.92mW.
It is clear that the central dip depth is reduced by the lower power density.
I've attached a photo of some changes to the cryocooler-tank connection design. We can save money and space by removing the 45 degree 1.33" conflat ports from the custom CH104 to 6" conflat adapter and using zero length conflat reducers at the unused 4 way cross ports, ie replace the 4-way piece blanks with holes for the vacuum line and gauge. The primary goal for these changes is to shorten the path from the cooler's heat station to the tank so that we keep the long thermal strap for use inside the tank. Also, the height is reduced from 89 cm to 59 cm.
A slightly different cupper adapter is needed to accomodate the thick strap, but no adapter will be needed anymore between the heat station and the thermal strap (same diameter round mates(new holes will need to be drilled though)).
Tightened all of the vacuum ports on the chamber so that the flange interfaces are all now metal-to-metal, ie full copper gasket compression. All of the ports required at least two star pattern passes before reaching this point, except for the bellows line to the turbo/backing pumps which was already at complete compression. Prior to tighening, the wide range gauge gave a pressure reading of 5.8x10-6 Torr and the ion-gauge showed 5.75x10-6 Torr. After tightening the wide range flange the reading dropped to 5.6x10-6 ; after tightening the ion-gauge flange the gauge reading dropped to 5.69x10-6.
For future reference: the 4.625" flanges use 5/16" torx bolts and 1/2" nuts, and the 2.75" flanges use a 1/4" torx bolt and 7/16" nuts.
Attached is a drawing of the first phase (minimal vibration isolation) cryocooler attachment, where the main tank connects via the blue rimmed feedthrough. Boxed/circled components are those that will require custom fabrication:
Currently there are only two connections that require viton o-ring rather than conflat connections (cooler to piece 1, piece 3 to HV feedthrough).
Clipping and saturation were investigated by the semi-analytical model. In the analysis, the waist radius of 20um at the micrometer position of 8mm is used.
Firstly, the clipping loss was just geometrically calculated. Here the saturation issue was completely ignored.ã€€The elements P6, P3, and P2 have the sizes of (500um)^2, (750um)^2m, and (1000um)^2, respectively. However, these numbers could not explain the clipping loss observed at the large spot sizes. Instead, empirically the effective sizes of (350um)^2, (610um)^2, and (860um)^2 were given to match the measurement and the calculation. This is equivalent to have 70um of an insensitive band at each edge of an element (Attachment 1). These effective element sizes are used for the calculation throughout this elog entry.
2) Saturation modeling
To incorporate the saturation effect, set a threshold power density. i.e. When the power density exceeds the threshold, the power density is truncated to this threshold. (Hard saturation)
Resulting loss was estimated using numerical integration using Mathematica. When the threshold power density was set to be 0.85W/mm^2, the drop of QE was approximately matched at the waist (Attachment 2). However, this did not explain the observed much-earlier saturation at the lower density. This suggests that the saturation is not such hard.
In order to estimate the threshold power density, look at the beam size where the first saturation starts. The earlier sagging of the QE was represented by the threshold density of 0.1W/mm^2. (Attachment 3)
InAsSb PD QE Test
The relationship between the spot radius and the apparent QE (EQE) was measured.
1) The spot size was checked with DataRay Beam'R2. The beam scanner was mounted on the post with a micrometer stage in the longitudinal direction. (Attachment1 upper plot)
It was confirmed that the beam is focused down to ~22um. The incident power was about 0.9mW.
2) The InAsSb detector (Sb3513A2) was mounted on the PD holder and then mounted on the stage+post. The photocurrent was amplified by a FEMTO's transimpedance amp (V/A=1e3Ohm). The dark current and the total photocurrent were measured at each measurement point with the beam aligned to the PD every time. The estimated EQEs were plotted in the lower plot of the attachment.
Note that P2, P3, and P6 elements have the size of (500um)^2, (750um)^2, and (1000um)^2, respectively.
The absolute longitudinal position of the sensor was of course slightly different from the position of the beam scanner. So the horizontal axis of the plots was arbitrary adjuted based on the symmetry.
The remarkable feature is that the QE goes down with small spot size. This is suggesting a nonlinear loss mechanism such as recombination loss when the carrier density is high.
With the present incident power, the beam size of 100um is optimal for all the element sizes. For the larger elements, a bigger beam size seems still fine.
The next step is to estimate the clipping loss and the saturation threshold with the Gaussian beam model.
The anti-alias boards in the QIL AA chassis have been replaced with newer ones I found in the EE shop (serial numbers S1200217, S1200274, S1200275, S1200277).
The new boards (D070081-v4) have input buffers and a reasonably high input impedance (20k), unlike the old boards (D070081-v1). However, according to the DCC revision notes, they may suffer from some excess low frequency noise caused by LT1492 opamps. If it becomes a problem for us, we can replace those opamps.
The low input impedance of the original boards explains the anomalous ADC/DAC loopback measurement Jon made several months ago. It should now be close to 0.5 ADC ct per DAC ct. I have checked the DC gain for the first few channels, but have not exhaustively tested the new boards. (Perhaps Jon has a script to automate this?)
From Cryo Cav setup
Borrowed ITC510 Laser Driver/TEC controller combo -> QIL
Ever since the initial pumpdown the pressure in the new cryo chamber has been stuck at ~6e-6 torr, so there's probably a small leak.
Our vacuum gauge controller has a serial communications port, which we can use to log the system pressure to aid in leak hunting. It's connected now to an unused port on fb4. A small python-based epics server queries the pressure gauges and makes the data available as two epics channels, C4:VAC-CRYO_PRES_P1 (wide range gauge) and C4:VAC-CRYO_PRES_P2 (ion gauge). These are recorded by the framebuilder. The script is stored under ~controls/services on fb4 and should start automatically on reboot.
A historical note - electricians from Facilities visited the lab several weeks ago and installed new electrical service. To do this with a minimum of disruption to the lab, they de-installed some electrical outlets along the south wall and reused the conductors. They also taped up plastic sheeting to the table enclosure to protect the squeezing and laser stabilization experiments.
The spectrum analyzer SR785 uses a Ethernet-Wifi converter GWU637 from IOGEAR to connect to the WIFI in the lab. Today I am trying to download experiment data from SR785 as always but somehow it cannot find the device anymore. After struggling for a while, I restart the GWU637 device (unplug and plug the power cable) and then I can download data again.
I think it needs to be restarted after running for a couple weeks.
we'll suffer if we use a oil based pump long term - please find and order a dry pump to back this turbo. Not only is it bad for the vac chamber, its bad for other optics in the lab,
We installed the oil filter/trap on the roughing pump and began pulling a vacuum. This was delayed due to the turbo pump flashing an error message and shutting off automaticallly after failing to spin up within its preset ramp-up time (error message 1221(3)). Upon restarting the system there were no ramp up issues, likely due to the chamber already having pumped down to ~0.5 Torr at the time of restart. Fix: need to increase the turbo pump delay start time, currently at 0 (immediately spins up). After 30 minutes of pumping the pressure reached ~4e-5 Torr in the chamber. (It reached 7e-6 torr after 3 days of pumping.)
We also tested an alternate internal configuration with the unwrapped components (see attached doodle, unlabelled green disc is the cold plate). This has the advantage of thermally isolating the outer radiation shield from the cold plate, but, we found, would slightly misalign the optical input ports.
Several ant traps were placed around the lab to combat the observed ant problem.
Today we unpacked the radiation shields and started to puzzle out how to assemble them. Attached are photos of the parts as we guessed they are intended to stack up. We didn't see how the outer shield would be supported and isolated from the cold plate, so we are contacting Rahul to clarify.
One detail not shown in these photos is the rather poor weld quality on the interior of the outer shield.
We closed the chamber without installing the radiation shields or cold plate in order to test the vacuum pressure of the empty system. Upon turning on the backing pump there was a fine oil mist from exhaust port, so the pump was turned off and an oil trap/filter has been purchased.
It was estimated that 12 inch/pound of torque is required for each of the top plate bolts in order to compress the 75 Shore hardness Viton o-ring by 20% (recommended by O'Hanlon).
Someone (not me) has recently changed the IP addresses of the lab machines. I see the new assignments are the following:
We tested the cryo cooler and the turbo pump this afternoon. We ran the cryo cooler for two hours. The equalization pressure is 275psi(the pressure before we turn it on) and operating pressure (pressure after running for two hours) is 295psi(attachment 1). The operating pressure is lower than expected; the manual indicates the pressure is expected to be 300-320psi.
We cycled the turbo pump. It appears to be functioning properly.
The lenses were arranged so that the spot on the PD can become smaller. A quick measurement on a (500um)^2 element showed the QE of ~80%
With the strong focusing lens of f=40mm, the beam was once expanded to a few mm. Then f=75mm lens focuses the beam to ~30um (radius). (See Attachments 1&2)
With this new beam, the QE was quickly checked. The new measurement is indicated as "Sb3513 A2P6new" in the plot. It showed the QE of ~80%.
The AOI was scanned to find any maximum, but the AOI of 0deg was the best at least with the given beam. I'm not sure yet why 500umx500um requires such small beam radius like 30um. Awesome
Fiber Collimator (Thorlabs F028APC-2000+AD11F+LMR1) and MIR sensor cards (Thorlabs VRC6S Qty2) were delivered.
The sensor card is liquid crystal and seems temperature sensitive. It's slow and diffused. But at least we can now see 2um beams in a certain condition.
The fiber collimator seems working fine, but this gave me another issue. Now because the beam is small (w<500um) everywhere, I can't focus it very well. To make a focused beam, one needs a large beam, of course. Previously, the beam was not well focused. Therefore the final focused beam with f=150mm was sufficiently small like w=50um.
It looks like some kind of telescope is necessary.